one number exceeds another number by 2 and the product of the two numbers is 143. Find the numbers
Answer by algebrapro18(206)

well create a system of equations to represent the situation

x+2=y
xy=143

now solve the system

xy=143
x(x+2)=143
x^2+2x=143
x^2+2x-143=0
(x+13)(x-11)=0
x+13=0
x = -13

x-11 = 0
x = 11

y=x+2
y=-13+2
y=-11

y=x+2
y=11+2
y=13